OpenStudy (anonymous):

Find an equation of the line having the given slope and containing the given point m=7/8,(2,-4)

OpenStudy (saifoo.khan):

y = mx + c -4 = 7/8(2) + c c = -23/4

OpenStudy (saifoo.khan):

y = mx + c y = 7 x -23 --- --- 8 4 8y = 7x - 46 ans